IP angles scan experiment - 05 March 2007

LER Single beam X-angle scan

Note: Total offset: angle with respect to the initial value

TimeStep (mu rad) Total Offset (mu rad)SMP Comments
14:45 . . . LER unstable: going to regular by-2 pattern.
14:46 . . . Moving PR04 X-Collimators by + 3mm
14:47 . . . LER still Longitudinally unstable. LFB not working well at 300 mA. Going to 350 mA.
14:58 . . . 300 nT bump in PR06. Moving collimators back to original value.
15:00 0 0 10 Start Scan
15:02 -300 -300 28 .
15:04 -300 -600 46 .
15:06 -300 -900 59 .
15:08 -300 -1200 71 .
15:11 -300 -1500 89 Large Bkg spikes in DCH, SVT.
15:15 -300 -1800 100 Problems with nu_X
15:22 0 -1800 118 No TI
15:23 +150 -1650 130 .
15:25 +300 -1350 150 .
15:27 +300 -1050 161 .
15:29 +300 -750 174 .
15:31 +300 -450 190 .
15:32 +300 -150 202
15:33 +300 +150 213 .
15:35 +300 +450 224 .
15:37 +300 +750 238 .
15:39 +300 +1050 253 Large Bkg
15:41 -150 +1000 267 .
15:43 -300 +700 282 .
15:45 -300 +400 300 .
15:49 -300 +100 315 .
15:50 -100 0 327 .